### v2

# The Six-Vertex Yang-Baxter Groupoid

Abstract.
A parametrized Yang-Baxter equation is a map from a group to a set of R-matrices, satisfying the Yang-Baxter commutation relation. For the six-vertex model, there are two main regimes of the Yang-Baxter equation: the free-fermionic point, and everything else. For the free-fermionic point, there exists a parametrized Yang-Baxter equation with a large parameter group $\operatorname{GL}(2)\times\operatorname{GL}(1)$. For non-free-fermionic six-vertex matrices, there are also parametrized Yang-Baxter equations, but these do not account for all possible interactions. Instead we will construct a groupoid parametrized Yang-Baxter equation that does reflect all possible Yang-Baxter equations in the six-vertex model.
1991 Mathematics Subject Classification:
16T25,17B38,81R12
1. Introduction
Let $V$ be a finite-dimensional vector space. If $r,s,t\in\operatorname{End}(V\otimes V)$, the Yang-Baxter commutator is defined to be the endomorphism of $V\otimes V\otimes V$ defined by
$\llbracket r,s,t\rrbracket=(r\otimes I_{V})(I_{V}\otimes s)(t\otimes I_{V})-(I_{V}\otimes t)(s\otimes I_{V})(I_{V}\otimes r).$
Let $G$ be a group. A (group) parametrized Yang-Baxter equation with parameter group $G$ is a map $\pi:G\longrightarrow\operatorname{End}(V\otimes V)$ such that
$\llbracket\pi(g),\pi(gh),\pi(g)\rrbracket=0.$
Parametrized Yang-Baxter equations are a mainstay of integrable lattice models, originating in Baxter’s method of proving the commutability of row transfer matrices in the six- and eight-vertex models.
The six-vertex model is perhaps the most basic example. Let $V=\mathbb{C}^{2}$, with standard basis $\mathbf{e}_{1},\mathbf{e}_{2}$. A six-vertex matrix is an endomorphism of $V\otimes V$, with basis $\mathbf{e}_{1}\otimes\mathbf{e}_{1},\mathbf{e}_{1}\otimes\mathbf{e}_{2}$, $\mathbf{e}_{2}\otimes\mathbf{e}_{1}$, $\mathbf{e}_{2}\otimes\mathbf{e}_{2}$, having a matrix of the form
(1) $u=\left(\begin{array}[]{cccc}a_{1}(u)&&&\\ &c_{1}(u)&b_{1}(u)&\\ &b_{2}(u)&c_{2}(u)&\\ &&&a_{2}(u)\end{array}\right).$
We will always consider only $u$ with $c_{1}(u),c_{2}(u)$ nonzero. The entries in the matrix may be referred to as Boltzmann weights due to the origin of the subject in statistical mechanics [1].
Let us give two examples of parametrized Yang-Baxter equations.
First, let $q_{1}$ and $q_{2}$ be fixed, nonzero complex numbers. For $z_{1},z_{2},w\in\mathbb{C}^{\times}$, define a six-vertex matrix $R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ to be the matrix with the following Boltzmann weights:
$\begin{array}[]{|l|l|l|l|l|l|}\hline\cr a_{1}(R)&a_{2}(R)&b_{1}(R)&b_{2}(R)&c_{1}(R)&c_{2}(R)\\ \hline\cr q_{1}z_{1}-q_{2}z_{2}&q_{1}z_{1}-q_{2}z_{2}&q_{1}(z_{1}-z_{2})&q_{2}(z_{1}-z_{2})&z_{1}w(q_{1}-q_{2})&z_{2}w^{-1}(q_{1}-q_{2})\\ \hline\cr\end{array}$
Then one may check that
$\llbracket R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w_{1}),R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1}z_{3},z_{2}z_{4},w_{1}w_{2}),R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{3},z_{4},w_{2})\rrbracket=0.$
So this is an example of a parametrized Yang-Baxter equation in the six-vertex model with parameter group (×$(\mathbb{C}^{\times})^{3}$.
A second example is, with $z_{1},z_{2},w\in\mathbb{C}^{\times}$, the $R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ to be the matrix with the following Boltzmann weights:
$\begin{array}[]{|l|l|l|l|l|l|}\hline\cr a_{1}(R)&a_{2}(R)&b_{1}(R)&b_{2}(R)&c_{1}(R)&c_{2}(R)\\ \hline\cr q_{1}z_{1}-q_{2}z_{2}&q_{1}z_{2}-q_{2}z_{1}&q_{1}(z_{1}-z_{2})&q_{2}(z_{1}-z_{2})&z_{1}w(q_{1}-q_{2})&z_{2}w^{-1}(q_{1}-q_{2})\\ \hline\cr\end{array}$
Again, we have
$\llbracket R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w_{1}),R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{1}z_{3},z_{2}z_{4},w_{1}w_{2}),R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{3},z_{4},w_{2})\rrbracket=0.$
So this is also a parametrized Yang-Baxter equation in the six-vertex model with parameter group (×$(\mathbb{C}^{\times})^{3}$.
These two examples are very similar: the weights are the same except for that $a_{2}$ entries, which differ. The weights $R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ are free-fermionic, meaning that they satisfy the condition
$a_{1}(R)a_{2}(R)+b_{1}(R)b_{2}(R)-c_{1}(R)c_{2}(R)=0.$
On the other hand the weights $R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ are not free fermionic, but instead satisfy the condition $a_{1}(R)=a_{2}(R)$.
Despite the similarity between the above two examples, there are nevertheless important differences between the free-fermionic regime and the non-free-fermionic regime in the six-vertex model. The free-fermionic example $R^{\operatorname{ff}}_{q_{1},q_{2}}$ can be extended to a parametrized Yang-Baxter equation with a larger group $\operatorname{GL}(2,\mathbb{C})\times\operatorname{GL}(1,\mathbb{C})$. In contrast $R^{\operatorname{cf}}_{q_{1},q_{2}}$ does not extend to a larger group.
As was shown by Drinfeld [5] and Jimbo [7], solutions to the Yang-Baxter equation are explained by the theory of quantum groups. For non-free-fermionic parametrized Yang-Baxter equations such as $R^{\operatorname{cf}}$, the relevant quantum group is $U_{q}(\mathfrak{g}\mathfrak{l}_{2})$ or its affinization, or Drinfeld twists, including two-parameter quantum groups. For free-fermionic Yang-Baxter equation such as $R^{\operatorname{ff}}$, the relevant quantum group is the superalgebra $U_{q}(\mathfrak{s}\mathfrak{l}(1|1))$, or its affinization ([2, 3]).
Now there are substantial differences between the representation theories of $U_{q}(\mathfrak{g}\mathfrak{l}_{2})$ and $U_{q}(\mathfrak{s}\mathfrak{l}(1|1))$. For example $U_{q}(\mathfrak{s}\mathfrak{l}(1|1))$ has many irreducible two-dimensional representations, accounting for the large $\operatorname{GL}(2)\times\operatorname{GL}(1)$-parametrized Yang-Baxter equation in the free-fermionic case. In contrast, this is not true for $U_{q}(\mathfrak{sl}_{2})$, which has only one. (Its affinization has a one-parameter family of two-dimensional representations.) Correspondingly, we do not find such a large (group) parametrized Yang-Baxter equation in the non-free-fermionic case.
Moreover, the group parametrized Yang-Baxter equations in the non-free-fermionic case do not account for all possible interactions between six-vertex matrices. To demonstrate this, let us consider the parametrized Yang-Baxter equation $R^{\operatorname{cf}}_{q_{1},q_{2}}$. We choose an element $r=R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ of this group. There are always six-vertex matrices $u$ and $v$ that are outside the group such that $\llbracket u,v,r\rrbracket=0$. Indeed, by Corollary 3.2 below, if $u$ is any six-vertex matrix such that $\Delta(u)=\Delta(r^{\ast})=(\frac{q_{1}+q_{2}}{q_{2}},\frac{q_{1}+q_{2}}{q_{1}})$, then there exists a $u$ such that $\llbracket u,v,r\rrbracket=0$. The condition on $u$ restricts it to a four-dimensional variety, but the group of matrices $R^{\operatorname{cf}}_{q_{1},q_{2}}$ is three-dimensional, so $u$ and $v$ that are outside the group certainly exist.
However if we expand the concept of a parametrized Yang-Baxter equation to allow the parameter object to be a groupoid, rather than a group, then we will show that all such interactions can be accounted for. Recall that a groupoid $\mathfrak{G}$ is a set with an associative composition law, which we will denote $\star$, that is only partially defined. Thus if $u,v\in\mathfrak{G}$, it may or may not be the case that $u,v$ is defined.
Thus we may define a groupoid parametrized Yang-Baxter equation to be a map $\pi:\mathfrak{G}\to\operatorname{End}(V\otimes V)$, where $\mathfrak{G}$ is a groupoid and $V$ a vector space, to be a map such that if $g,h\in\mathfrak{G}$ are such that $g\star h$ is defined, then $\llbracket\pi(g),\pi(g\star h),\pi(h)\rrbracket=0$. We will construct such a groupoid parametrized Yang-Baxter equation. The groupoid $\mathfrak{G}$ will decompose into a disjoint union of a free-fermionic part $\mathfrak{G}_{\operatorname{ff}}$, which is actually the group $\operatorname{GL}(2)\times\operatorname{GL}(1)$, and a more interesting non-free-fermionic part $\mathfrak{G}_{\operatorname{nf}}$, which is a groupoid.
This groupoid parametrized Yang-Baxter equation has the stronger property that $g\star h$ is defined if and only if $\llbracket\pi(g),w,\pi(h)\rrbracket=0$ has a solution, and if this is the case then $w$ is a constant multiple of $\pi(g\star h)$. This strong property justifies our assertion that the groupoid parametrized Yang-Baxter equation accounts for all interactions in the six-vertex model.
Turning to the non-free-fermionic groupoid, with $V=\mathbb{C}^{2}$, the map $\pi:\mathfrak{G}_{\operatorname{nf}}\to\operatorname{End}(V\otimes V)$ is not an isomorphism onto its image. The set $\pi(\mathfrak{G}_{\operatorname{nf}})$ is a specific set $\overline{\Omega}$ of matrices in $\operatorname{End}(V\otimes V)$. The set $\overline{\Omega}$ can be thought of as a quasi-affine variety, and it has a Zariski topology. The set $\pi:\mathfrak{G}_{\operatorname{nf}}\rightarrow\overline{\Omega}$ is a morphism that is a birational equivalence. Essentially, $\mathfrak{G}_{\operatorname{nf}}$ is obtained by “blowing up” certain lower-dimensional boundary components in a specific way. We refer to the last section for details.
The set $\overline{\Omega}$ has an open subset $\Omega^{\circ}$, and three boundary components. A special role is played by two functions on $\overline{\Omega}$. These are:
$\Delta_{1}(u)=\frac{a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)}{a_{1}(u)b_{1}(u)},$
$\Delta_{2}(u)=\frac{a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)}{a_{2}(u)b_{2}(u)}.$
Let $\Delta(u)=(\Delta_{1}(u),\Delta_{2}(u))$. This is a function on $\Omega^{\circ}$ that is not defined on certain boundary components, but it can be defined on the blowup $\mathfrak{G}_{\operatorname{nf}}$. Now if $u,v\in\mathfrak{G}_{\operatorname{nf}}$, then $u\star v$ is defined if and only if $\Delta(u)=\Delta(v^{\prime})$, and if $w=u\star v$ then $\Delta(w)=\Delta(v)$.
Remark 1.
Define $\Delta_{0}(u)=\Delta_{1}(u)\Delta_{2}(u)$. If $\delta_{0}$ is given, let
$\mathfrak{G}_{\operatorname{nf}}(d_{0})=\{u\in\mathfrak{G}_{\operatorname{nf}}|\Delta_{0}(u)=\delta_{0}\}.$
Then we have a further decomposition
$\mathfrak{G}_{\operatorname{nf}}=\bigsqcup_{d_{0}}\mathfrak{G}_{\operatorname{nf}}(d_{0})$
Indeed, Proposition 3.4 shows that if $\llbracket u,w,v\rrbracket=0$ then $\Delta_{0}(u)=\Delta_{0}(w)=\Delta_{0}(v)$. This further decomposition of the groupoid is certainly important but has no role in the proofs.
One may speculate that there are other Yang-Baxter equations parametrized by groupoids. Another paper relating parametrized Yang-Baxter equations to groupoids is Felder and Ren [6]. However their use of a groupoid is different from ours.
This paper is based on [9], to which it adds analysis of the boundary components of the Six-Vertex Groupoid.
We thank Amol Aggarwal, Andrew Hardt and Travis Scrimshaw for helpful conversations about the groupoid.
2. Groupoids
A groupoid is a set $G$ with a partially defined composition. This consists of a map $\mu:S\longrightarrow G$, where $S$ is a subset of $G\times G$. If $a,b\in G$ we say that the product $a\star b$ is defined if $(a,b)\in S$, and then we write $a\star b=\mu(a,b)$. The groupoid is also required to have an “inverse map” $x\mapsto x^{\prime}$ from $G\rightarrow G$. The inverse map is more commonly denoted as $x\mapsto x^{-1}$, but we will be concerned with a groupoid whose elements are matrices, and we will reserve the notation $x^{-1}$ for the matrix inverse. The following axioms are required.
Axiom 1 (Associative Law).
If $a\star b$ and $b\star c$ are defined then $(a\star b)\star c$ and $a\star(b\star c)$ are defined, and they are equal.
We say that $a\star b\star c$ is defined if $a\star b$ and $b\star c$ are defined, and then we denote $(a\star b)\star c=a\star(b\star c)$ as $a\star b\star c$.
Axiom 2 (Inverse).
The compositions $a\star a^{\prime}$ and $a^{\prime}\star a$ are always defined. Thus if $a\star b$ is defined, then $a\star b\star b^{\prime}$ is defined, and this is required to equal $a$. Similarly $a^{\prime}\star a\star b$ is defined, and this is required to equal $b$.
Example 2.1.
A category $\mathcal{C}$ is small if its class of objects is a set. A small category is a groupoid category if every morphism is an isomorphism. Assuming this, the disjoint union
$G=\bigsqcup_{A,B\in\mathcal{C}}\mathrm{Hom}(A,B)$
is a groupoid, with the $\star$ operation being composition: thus if $a\in\mathrm{Hom}(A,B)$ and $b\in\mathrm{Hom}(C,D)$, then $a\star b$ is defined if and only if $B=C$. The groupoid axioms are clear.
Lemma 2.2.
In a groupoid, we have $(a^{\prime})^{\prime}=a$. Moreover if $a\star b$ is defined then so is $b^{\prime}\star a^{\prime}$ and $(a\star b)^{\prime}=b^{\prime}\star a^{\prime}$.
Proof.
Since $(a^{\prime})^{\prime}\star a^{\prime}$ and $a^{\prime}\star a$ are both defined, by the Associative Law the product $(a^{\prime})^{\prime}\star a^{\prime}\star a$ is defined, and using the Inverse Axiom, this equals both $(a^{\prime})^{\prime}$ and $a$. For the second assertion, assume $a\star b$ is defined. It follows from the axioms that
$(a\star b)^{\prime}=(a\star b)^{\prime}\star a\star b\star b^{\prime}\star a^{\prime}=b^{\prime}\star a^{\prime}.\qed$
Given a groupoid $G$, let us say an element $A$ is idempotent if $A\star A$ is defined and $A\star A=A$.
Lemma 2.3.
An element $A\in G$ is an idempotent if and only if $A=g\star g^{\prime}$ for some $g\in G$. If $A$ is idempotent then $A=A^{\prime}$.
Proof.
It is easy to check that $g\star g^{\prime}$ is idempotent. Conversely if $A$ is idempotent, then $A=A\star A^{\prime}$ since $A=A\star A=A\star A\star A^{\prime}=A\star A^{\prime}$, and so $A$ can be written $g\star g^{\prime}$ with $g=A$. Now if $A=g\star g^{\prime}$ then $A=A^{\prime}$ as a consequence of Lemma 2.2. ∎
Lemma 2.4.
If $g\in G$ then there are unique idempotents $A$ and $B$ such that $g=g\star A$ and $g=B\star g$.
Proof.
We can take $A=g^{\prime}\star g$, and this is an idempotent such that $g\star A=g$. Conversely if $A^{\prime}$ is any other element such that $g\star A^{\prime}=g$, then $g^{-1}\star g=g^{-1}\star g\star A^{\prime}=A^{\prime}$, so $A^{\prime}=A$. The statements about $B$ are proved similarly. ∎
Proposition 2.5.
Let $G$ be a groupoid. Then there exists a groupoid category whose groupoid (as in Example 2.1) is isomorphic to $G$.
Proof.
Let us define a category $\mathcal{C}$ whose objects and morphisms are all elements of $G$. The objects are the idempotent elements of $G$. If $A,B$ are objects, we define $\mathrm{Hom}(A,B)$ to be the set of $g\in G$ such that $A\star g=g$ and $g\star B=B$. By Lemma 2.4,
$G=\bigsqcup_{A,B\in\mathcal{C}}\mathrm{Hom}(A,B).$
We must show that if $f\in\mathrm{Hom}(A,B)$ and $g\in\mathrm{Hom}(B,C)$ then $g\star f$ is defined and is in $\mathrm{Hom}(A,B)$. We can write $f=B\star f$ and $g=g\star B$, and then $g\star f=g\star B\star B\star f$ is defined since $B\star B$ is defined. It is clear that $C\star g\star f=g\star f$ and $g\star f\star A=g\star f$, so $g\star f\in\mathrm{Hom}(A,C)$.
Now an idempotent $A$ is itself both an object of the category and a morphism in $\mathrm{Hom}(A,A)$; to distinguish this double role we denote it as $1_{A}$ when regarding it as a morphism. The category axioms are easily checked. ∎
If $G_{i}$ ($i\in I$) is a parametrized family of groupoids, then the disjoint union $G=\sqcup G_{i}$ is naturally a groupoid. For example a disjoint union of groups is a groupoid.
Lemma 2.6.
If $G$ is a groupoid, and if $A$ is an idempotent, then
$\mathrm{Aut}(A):=\{g\in G|g\star A=A\star g=g\}$
is a group.
Proof.
Let us check that $\mathrm{Aut}(A)$ is closed under $\star$. If $g,h\in\mathrm{Aut}(A)$, then $g\star A$ and $A\star h$ are defined so $g\star A\star h$ is defined. This equals $g\star A\star A\star h=g\star h$. We leave the remaining details to the reader. ∎
3. Yang-Baxter equation for the six-vertex model
The six-vertex model in statistical mechanics can be described algebraically in terms of the matrices of weights for each vertex. See [4], Section 1 for details. We study the Yang-Baxter equation for the matrices which arise from the six-vertex model.
Definition 1.
A six-vertex matrix $u\in\operatorname{GL}_{4}(\mathbb{C})$ is a $4\times 4$ matrix of the form (1) where $c_{1}(u)$ and $c_{2}(u)$ are nonzero. Let $S$ be the set of the six-vertex matrices. Let $S^{\times}$ be the set of invertible elements of $S$. Also let $S^{\bullet}$ be the subset of $S$ in which all six coefficients $a_{i}(u)$, $b_{i}(u)$, $c_{i}(u)$ are nonzero. Let $S^{\circ}=S^{\times}\cap S^{\bullet}$ be the subset of elements of $S^{\bullet}$ that are invertible, so that furthermore $c_{1}(u)c_{2}(u)-b_{1}(u)b_{2}(u)\neq 0$.
We will require a number of different subsets of $S$. These are all locally closed in either the Zariski or complex topologies. These are summarized in Table 1.
$S$ $c_{1},c_{2}\neq 0$
$S^{\times}$ $a_{1},a_{2},c_{1},c_{2},c_{1}c_{2}-b_{1}b_{2}\neq 0$
$S^{\bullet}$ $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2}\neq 0$
$S^{\circ}$ $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2},c_{1}c_{2}-b_{1}b_{2}\neq 0$
$S_{\operatorname{ff}}$ $c_{1},c_{2}\neq 0$, $N(u)=0$
$\Omega$ $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2},N(u)\neq 0$;  $(\Omega=\Omega^{\circ}\cup\Omega_{B})$
$\Omega^{\circ}$ $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2},c_{1}c_{2}-b_{1}b_{2},N(u)\neq 0$
$\Omega_{b}$ $a_{1},a_{2},c_{1},c_{2}\neq 0,b_{1}=b_{2}=0,a_{1}a_{2}-c_{1}c_{2}=0$
$\Omega_{a}$ $b_{1},b_{2},c_{1},c_{2}\neq 0,a_{1}=a_{2}=0,b_{1}b_{2}-c_{1}c_{2}=0$
$\Omega_{B}$ $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2},N(u)\neq 0,c_{1}c_{2}-b_{1}b_{2}=0$
$\overline{\Omega}$ $\Omega^{\circ}\cup\Omega_{b}\cup\Omega_{a}\cup\Omega_{B}$
Table 1. Various subsets of the six-vertices $u\in S$, including the constraints that define them. All sets require $c_{1}=c_{1}(u)$ and $c_{2}=c_{2}(u)$ to be nonzero. The sets labeled $\Omega$ (with various decorations) are needed for the analysis of the non-free-fermionic part of the Yang-Baxter groupoid, although $\Omega_{a}$ and $\Omega_{b}$ are free-fermionic.
We will denote by $B(u)=\left(\begin{smallmatrix}c_{1}(u)&b_{1}(u)\\ b_{2}(u)&c_{2}(u)\end{smallmatrix}\right)$ the middle $2\times 2$ block. We will also denote
$N(u)=a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)=a_{1}(u)a_{2}(u)-\det(B(u)).$
Let $V\cong\mathbb{C}^{2}$ with the standard basis $e_{1},e_{2}$. We view a six-vertex matrix as a matrix of an operator $u\in\operatorname{End}(V\otimes V)$ in basis $e_{1}\otimes e_{1},e_{1}\otimes e_{2},e_{2}\otimes e_{1},e_{2}\otimes e_{2}$. By abuse of notation, we denote by $u$ both the six-vertex matrix and the corresponding operator.
If $a_{1}(u)$ and $a_{2}(u)$ are nonzero we define
$u^{\ast}=\left(\begin{array}[]{cccc}a_{1}(u^{\ast})&&&\\ &c_{1}^{\ast}(u)&b_{1}^{\ast}(u)&\\ &b_{2}^{\ast}(u)&c_{2}^{\ast}(u)&\\ &&&a_{2}^{\ast}(u)\end{array}\right)=\left(\begin{array}[]{cccc}a_{1}(u^{\ast})&&&\\ &c_{2}(u)&-b_{1}(u)&\\ &-b_{2}(u)&c_{1}(u)&\\ &&&a_{2}^{\ast}(u)\end{array}\right),$
where $a_{1}(u^{\ast})$ and $a_{2}(u^{\ast})$ are defined by
$a_{1}(u^{\ast})=\frac{c_{1}(u)c_{2}(u)-b_{1}(u)b_{2}(u)}{a_{1}(u)},\qquad a_{1}(u^{\ast})=\frac{c_{1}(u)c_{2}(u)-b_{1}(u)b_{2}(u)}{a_{2}(u)}.$
If furthermore $u$ is invertible (or equivalently if $\det(B(u))\neq 0$) then $u^{\ast}=\det(B(u))u^{-1}$. A computation shows that
(2) $N(u^{\ast})=-\frac{\det(B(u))}{a_{1}(u)a_{2}(u)}N(u).$
If $u\in S^{\bullet}$ define
(3) $\Delta_{1}(u)=\frac{N(u)}{a_{1}(u)b_{1}(u)}=\frac{a_{2}(u)-a_{1}^{\ast}(u)}{b_{1}(u)},$
(4) $\Delta_{2}(u)=\frac{N(u)}{a_{2}(u)b_{2}(u)}=\frac{a_{1}(u)-a_{2}^{\ast}(u)}{b_{2}(u)}.$
Also define $\Delta:S^{\bullet}\longrightarrow\mathbb{C}^{2}$ by $\Delta(u)=(\Delta_{1}(u),\Delta_{2}(u))$. If $u\in S^{\circ}$, then $u^{\ast}\in S^{\bullet}$, so $\Delta(u^{\ast})$ is defined, and it follows from (2) that
(5) $\Delta_{1}(u^{\ast})=\frac{N(u)}{a_{2}(u)b_{1}(u)},\qquad\Delta_{2}(u^{\ast})=\frac{N(u)}{a_{1}(u)b_{2}(u)}.$
The functions $\Delta_{1}(u),\Delta_{2}(u),\Delta_{1}(u^{\ast})$ and $\Delta_{2}(u^{\ast})$ are thus regular on $S^{\bullet}$.
Let $V\cong\mathbb{C}^{2}$. For $u,w,v\in\operatorname{End}(V\otimes V)$, define the Yang-Baxter commutator on $V\otimes V\otimes V$:
(6) $\llbracket u,w,v\rrbracket=(u\otimes 1)(1\otimes w)(v\otimes 1)-(1\otimes v)(w\otimes 1)(1\otimes u).$
Then the Yang-Baxter equation is the identity
(7) $\llbracket u,w,v\rrbracket=0,\quad u,w,v\in\operatorname{End}(V\otimes V).$
We call a solution $(u,w,v)$ to (7) normalized if
(8) $c_{1}(w)=c_{1}(u)\,c_{1}(v),\qquad c_{2}(w)=c_{2}(u)\,c_{2}(v).$
Theorem 3.1.
Assume that $u,v$ are six-vertex matrices with $a_{1}(u),a_{2}(u),a_{1}(v),a_{2}(v)$ all nonzero so that $u^{\ast}$ and $v^{\ast}$ are defined. A necessary and sufficient condition that there exists a six-vertex matrix $w$ with $\llbracket u,w,v\rrbracket=0$ is that:
(9) $\frac{N(u)b_{1}(v)}{a_{1}(u)}=\frac{N(v)b_{1}(u)}{a_{2}(v)},\qquad\qquad\frac{N(u)b_{2}(v)}{a_{2}(u)}=\frac{N(v)b_{2}(u)}{a_{1}(v)}.$
The solution $w$ is unique up to scalar multiple, and there is a unique normalized solution determined by (8) and
(10) $\begin{array}[]{l}a_{1}(w)=a_{1}(u)a_{1}(v)-b_{2}(u)b_{1}(v),\qquad a_{2}(w)=a_{2}(u)a_{2}(v)-b_{1}(u)b_{2}(v),\\ b_{1}(w)=a_{1}(u^{\ast})b_{1}(v)+b_{1}(u)a_{1}(v)=b_{1}(u)a_{2}(v^{\ast})+a_{2}(u)b_{1}(v),\\ b_{2}(w)=a_{2}(u^{\ast})b_{2}(v)+b_{2}(u)a_{2}(v)=b_{2}(u)a_{1}(v^{\ast})+a_{1}(u)b_{2}(v).\end{array}$
The equivalence of the alternative expressions for $b_{1}(w)$ and $b_{2}(w)$ follows from (9).
Proof.
There are 14 equations that must be satisfied for the matrix $\llbracket u,w,v\rrbracket$ to vanish. However one equation is duplicated. The duplicated equation is
(11) $c_{1}(u)c_{2}(w)c_{1}(v)=c_{2}(u)c_{1}(w)c_{2}(v)$
The remaining twelve equations are:
(12) $\begin{array}[]{rcl}c_{1}(u)a_{1}(w)b_{2}(v)+b_{2}(u)c_{1}(w)c_{2}(v)&=&c_{1}(u)b_{2}(w)a_{1}(v)\\ b_{2}(u)c_{1}(w)b_{1}(v)+c_{1}(u)a_{1}(w)c_{1}(v)&=&a_{1}(u)c_{1}(w)a_{1}(v)\\ c_{2}(u)a_{1}(w)b_{2}(v)+b_{2}(u)c_{2}(w)c_{1}(v)&=&c_{2}(u)b_{2}(w)a_{1}(v)\\ c_{2}(u)c_{1}(w)b_{1}(v)+b_{1}(u)a_{1}(w)c_{1}(v)&=&a_{1}(u)b_{1}(w)c_{1}(v)\\ c_{2}(u)c_{1}(w)b_{2}(v)+b_{2}(u)a_{2}(w)c_{1}(v)&=&a_{2}(u)b_{2}(w)c_{1}(v)\\ b_{1}(u)c_{1}(w)b_{2}(v)+c_{1}(u)a_{2}(w)c_{1}(v)&=&a_{2}(u)c_{1}(w)a_{2}(v)\\ b_{2}(u)c_{2}(w)b_{1}(v)+c_{2}(u)a_{1}(w)c_{2}(v)&=&a_{1}(u)c_{2}(w)a_{1}(v)\\ c_{1}(u)c_{2}(w)b_{1}(v)+b_{1}(u)a_{1}(w)c_{2}(v)&=&a_{1}(u)b_{1}(w)c_{2}(v)\\ c_{1}(u)c_{2}(w)b_{2}(v)+b_{2}(u)a_{2}(w)c_{2}(v)&=&a_{2}(u)b_{2}(w)c_{2}(v)\\ c_{1}(u)a_{2}(w)b_{1}(v)+b_{1}(u)c_{1}(w)c_{2}(v)&=&c_{1}(u)b_{1}(w)a_{2}(v)\\ b_{1}(u)c_{2}(w)b_{2}(v)+c_{2}(u)a_{2}(w)c_{2}(v)&=&a_{2}(u)c_{2}(w)a_{2}(v)\\ c_{2}(u)a_{2}(w)b_{1}(v)+b_{1}(u)c_{2}(w)c_{1}(v)&=&c_{2}(u)b_{1}(w)a_{2}(v)\end{array}$
Equation (11) implies that if a solution $w$ exists, a unique constant multiple is normalized. Therefore we may impose the condition (8). Substituting these values for $c_{1}(w)$ and $c_{2}(w)$, each of the twelve equations is now divisible by one of $c_{1}(u),c_{2}(u),c_{1}(v)$ or $c_{2}(v)$, and on dividing these away, each of the twelve equations occurs twice and there are only six unique equations to be satisfied. These are
$\begin{array}[]{rcc}b_{2}(u)c_{1}(v)c_{2}(v)-b_{2}(w)a_{1}(v)+a_{1}(w)b_{2}(v)&=&0\\ -a_{1}(u)a_{1}(v)+b_{2}(u)b_{1}(v)+a_{1}(w)&=&0\\ c_{1}(u)c_{2}(u)b_{1}(v)+b_{1}(u)a_{1}(w)-a_{1}(u)b_{1}(w)&=&0\\ c_{1}(u)c_{2}(u)b_{2}(v)+b_{2}(u)a_{2}(w)-a_{2}(u)b_{2}(w)&=&0\\ -a_{2}(u)a_{2}(v)+b_{1}(u)b_{2}(v)+a_{2}(w)&=&0\\ b_{1}(u)c_{1}(v)c_{2}(v)-b_{1}(w)a_{2}(v)+a_{2}(w)b_{1}(v)&=&0\end{array}$
From the second and fifth equation we see that $a_{1}(w)$ and $a_{2}(w)$ must have the values in (10). Substituting these values in the remaining four equations we obtain
$\begin{array}[]{rcc}a_{1}(u)a_{1}(v)b_{2}(v)-b_{2}(u)b_{1}(v)b_{2}(v)+b_{2}(u)c_{1}(v)c_{2}(v)-b_{2}(w)a_{1}(v)&=&0\\ a_{1}(u)b_{1}(u)a_{1}(v)-b_{1}(u)b_{2}(u)b_{1}(v)+c_{1}(u)c_{2}(u)b_{1}(v)-a_{1}(u)b_{1}(w)&=&0\\ a_{2}(u)b_{2}(u)a_{2}(v)-b_{1}(u)b_{2}(u)b_{2}(v)+c_{1}(u)c_{2}(u)b_{2}(v)-a_{2}(u)b_{2}(w)&=&0\\ a_{2}(u)a_{2}(v)b_{1}(v)-b_{1}(u)b_{1}(v)b_{2}(v)+b_{1}(u)c_{1}(v)c_{2}(v)-b_{1}(w)a_{2}(v)&=&0\end{array}$
The second and fourth are each equivalent one of the two expressions for $b_{1}(w)$ in (10), and similarly the first and third equations are each equivalent to one of the two expressions for $b_{2}(w)$. The two expressions for $b_{1}(w)$ are equivalent if and only if (9) is satisfied, and similarly for $b_{2}(w)$. We have proved that a solution exists if and only if (9) is satisfied, and if so, there is a unique normalized solution. ∎
Remark 2.
Conditions (9) can also be written
(13) $\begin{array}[]{c}(a_{1}(u^{\ast})-a_{2}(u))b_{1}(v)=(a_{2}(v^{\ast})-a_{1}(v))b_{1}(u)=0,\\ (a_{2}(u^{\ast})-a_{1}(u))b_{2}(v)=(a_{1}(v^{\ast})-a_{2}(v))b_{2}(u)=0.\end{array}$
Corollary 3.2.
If $u,v\in S^{\bullet}$ a necessary and sufficient condition that there exists $w\in S$ such that $\llbracket u,w,v\rrbracket=0$ is that $\Delta(u)=\Delta(v^{\ast})$. If this is satisfied, then $w$ is determined up to constant multiple, and may be normalized as in (10).
Proof.
In Theorem 3.1, we only assumed that $a_{1}(u)$, $a_{2}(u)$, $a_{1}(v)$ and $a_{2}(v)$ are nonzero. If $u,v\in S^{\bullet}$, that is, if $b_{1}(u),b_{2}(u),b_{1}(v)$ and $b_{2}(v)$ are nonzero, then the two equations in (9) are equivalent to $\Delta_{1}(u)=\Delta_{1}(v^{\ast})$ and $\Delta_{2}(u)=\Delta_{2}(u^{\ast})$. ∎
Observe that if $u,v\in S^{\bullet}$ satisfy $\Delta(u)=\Delta(v^{\ast})$, then Theorem 3.1 guarantees that there is $w\in S$ such that $\llbracket u,w,v\rrbracket=0$, but it does not guarantee that $w\in S^{\bullet}$. Also $S^{\bullet}$ is not closed under the map $u\mapsto u^{\ast}$. However the set $S^{\circ}$ of invertible elements in $S^{\bullet}$ is closed under $u\mapsto u^{\ast}$. The set $S^{\circ}$ is open in $S$, so conclusions we draw in this case hold “generically.”
Lemma 3.3.
Suppose that $u,v,w\in S^{\times}$ satisfy $\llbracket u,w,v\rrbracket=0$. Then
$\llbracket u,w,v\rrbracket=\llbracket u^{\ast},v,w\rrbracket=\llbracket w,u,v^{\ast}\rrbracket=\llbracket v,u^{\ast},w^{\ast}\rrbracket=\llbracket w^{\ast},v^{\ast},u\rrbracket=\llbracket v^{\ast},w^{\ast},u^{\ast}\rrbracket=0.$
Proof.
Note that $u,v$ and $w$ are invertible $S^{\times}$. In the identity
$(u\otimes I)(I\otimes w)(v\otimes I)-(I\otimes v)(w\otimes I)(I\otimes u)=\llbracket u,w,v\rrbracket=0,$
multiplying on the left by $u^{-1}\otimes I$ and on the right by $I\otimes u^{-1}$ gives $\llbracket u^{-1},w,v\rrbracket=0$. Then multiplying by $\det(B(u))$ gives $\llbracket u^{\ast},w,v\rrbracket=0$. The identity $\llbracket w,u,v^{\ast}\rrbracket=0$ is proved similarly. Applying the operations $(u,w,v)\mapsto(u^{\ast},v,w)$ and $(u,w,v)\mapsto(w,u,v^{\ast})$ gives the six identities. ∎
In the following Proposition, $\Delta_{0}(u):=\Delta_{1}(u)\Delta_{2}(u)$. It follows by comparing (3) and (4) with (5) that
(14) $\Delta_{0}(u)=\Delta_{0}(u^{\ast}).$
Proposition 3.4.
If $u,w,v\in S^{\circ}$ satisfy $\llbracket u,w,v\rrbracket=0$, then
$\Delta(u)=\Delta(w),\qquad\Delta(u)=\Delta(v^{\ast}),\qquad\Delta(w^{\ast})=\Delta(u^{\ast}).$
Moreover
$\Delta_{0}(u)=\Delta_{0}(v)=\Delta_{0}(w).$
Proof.
The first assertion follows from Theorem 3.1 and Lemma 3.3. The second fact follow from (14). ∎
If $u\in S$ and $N(u)=0$, the vertex $u$ is called free-fermionic. Let $S_{\operatorname{ff}}$ be the set of all free-fermionic matrices. Observe that if $u\in S_{\operatorname{ff}}$ then $a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)$ is automatically nonzero, since it equals $c_{1}(u)c_{2}(u)\neq 0$.
Lemma 3.5.
If $u$ is free-fermionic and $a_{1}(u),a_{2}(u)$ are nonzero, so that $u^{\ast}$ is defined, then
(15) $a_{1}(u^{\ast})=a_{2}(u),\qquad a_{2}(u^{\ast})=a_{1}(u).$
The map $u\mapsto u^{\ast}$ extends to a continuous map on all free-fermionic elements of $S$.
Proof.
If $a_{1}(u)$ and $a_{2}(u)$ are nonzero, then (15) follows from $c_{1}(u)c_{2}(u)-b_{1}(u)b_{2}(u)=a_{1}(u)a_{2}(u)$. This implies (15). Using this formula for $a_{1}(u^{\ast})$ and $a_{2}(u^{\ast})$ gives the continuous extension to all free-fermionic six-vertex matrices. ∎
4. The Free-Fermionic Yang-Baxter equation
If $u,v$ are free-fermionic, then Theorem 3.1 simplifies as follows. Versions of this were proved by Korepin (see [8], page 126), and by Brubaker, Bump and Friedberg [4].
Theorem 4.1 (Korepin; Brubaker-Bump-Friedberg).
Suppose $u,v\in S$ are free-fermionic. Then there exists $w\in S$ that is also free-fermionic, such that $\llbracket u,w,v\rrbracket=0$. We have
(16) $\begin{array}[]{lc}a_{1}(w)=a_{1}(u)a_{1}(v)-b_{2}(u)b_{1}(v),&a_{2}(w)=-b_{1}(u)b_{2}(v)+a_{2}(u)a_{2}(v),\\ b_{1}(w)=b_{1}(u)a_{1}(v)+a_{2}(u)b_{1}(v),&b_{2}(w)=a_{1}(u)b_{2}(v)+b_{2}(u)a_{2}(v)\end{array}$
Proof.
First assume that $u,v\in S^{\circ}$. When $a_{1}(u),a_{2}(u)\neq 0$, the free-fermionic condition is equivalent to $\Delta(u)=(0,0)$. Thus by Corollary 3.2, condition (9) is satisfied and by Theorem 3.1 there exists $w\in S$ such that $\llbracket u,w,v\rrbracket=0$. Equation (16) follows from (10) taking (15) into account. The alternative expressions for $b_{1}(w)$ and $b_{2}(w)$ reduce to the same formula in this case. We see that there exists an open subset $U$ of $S_{\operatorname{ff}}\times S_{\operatorname{ff}}$ such that if $(u,v)\in U$, then $w$ is in the dense open subset $S^{\circ}$ of $S$. In this case $\Delta(w)=\Delta(v)=(0,0)$ by Propostion 3.4, so $w$ is free-fermionic. These arguments rely on the assumption that $(u,v)$ lies in an open subset of $S_{\operatorname{ff}}\times S_{\operatorname{ff}}$ but since the right-hand side of (16) is obviously continuous on all $S_{\operatorname{ff}}\times S_{\operatorname{ff}}$, the general case follows by continuity. ∎
$\left(\begin{array}[]{cc}a_{1}(u)&-b_{2}(u)\\ b_{1}(u)&a_{2}(u)\end{array}\right)\left(\begin{array}[]{cc}a_{1}(v)&-b_{2}(v)\\ b_{1}(v)&a_{2}(v)\end{array}\right)=\left(\begin{array}[]{cc}a_{1}(w)&-b_{2}(w)\\ b_{1}(w)&a_{2}(w)\end{array}\right)$
This result may be explained in terms of a parametrized Yang-Baxter equation. Let $\mathfrak{G}_{\operatorname{ff}}$ be the group $\operatorname{GL}(2,\mathbb{C})\times\mathbb{C}^{\times}$ and let $R_{\operatorname{ff}}:\mathfrak{G}_{\operatorname{ff}}\longrightarrow S_{\operatorname{ff}}$ be the bijective map
$R_{\operatorname{ff}}\left(\left(\begin{array}[]{cc}a_{1}&-b_{2}\\ b_{1}&a_{2}\end{array}\right),c_{1}\right)=\left(\begin{array}[]{cccc}a_{1}&&&\\ &c_{1}&b_{2}&\\ &b_{1}&c_{2}&\\ &&&a_{2}\end{array}\right),\qquad c_{2}=\frac{a_{1}a_{2}+b_{1}b_{2}}{c_{1}}.$
Corollary 4.2.
The map $R_{\operatorname{ff}}:\mathfrak{G}_{\operatorname{ff}}\longrightarrow S_{\operatorname{ff}}$ is a parametrized Yang-Baxter equation with parameter group $\Gamma_{\operatorname{ff}}\cong\operatorname{GL}(2,\mathbb{C})\times\operatorname{GL}(1,\mathbb{C})$. Thus if $\gamma,\delta\in\mathfrak{G}_{\operatorname{ff}}$ then
$\llbracket R_{\operatorname{ff}}(\gamma),R_{\operatorname{ff}}(\gamma\delta),R_{\operatorname{ff}}(\delta)\rrbracket=0.$
Proof.
Indeed (16) implies that $w=R_{\operatorname{ff}}(\gamma\delta)$, where $\gamma\delta$ is just the product of $\gamma$ and $\delta$ in the group. ∎
5. Preparations for the Non-Free-Fermionic Groupoid
We now turn to the non-free-fermionic case. In order to obtain a groupoid parametrized Yang-Baxter equation we will have to “blow up” part of the boundary of the free-fermionic domain. We will carry out these details in the next section, but here we make some preparations.
We have introduced several sets $\Omega^{\circ},\Omega_{b},\Omega_{a}$ and $\Omega_{B}$ in Table 1, of which $\Omega^{\circ}$ is the open subset of $S$ defined by the nonvanishing of $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2}$, $\det(B(u))$ and $N(u)$. We refer to the table for the definitions of the other three sets. The set $\Omega_{B}$ is of codimension 1 in $S$, and the subsets $\Omega_{b}$ and $\Omega_{a}$ are smaller, of codimension 3. The sets $\Omega_{b}$ and $\Omega_{a}$ are free-fermionic. Elements of $\Omega^{\circ}$ and $\Omega_{B}$ are not. Elements of $\Omega^{\circ}$ and $\Omega_{b}$ are invertible, but elements of $\Omega_{a}$ and $\Omega_{B}$ are not. We let $\overline{\Omega}$ be the union of the four disjoint sets $\Omega^{\circ},\Omega_{b},\Omega_{a}$ and $\Omega_{B}$. We will also denote $\partial\Omega=\Omega_{b}\cup\Omega_{a}\cup\Omega_{B}$. We will describe $\Omega^{\circ}$ as the interior of $\overline{\Omega}$ and $\partial\Omega$ as the boundary. Also let $\Omega=\Omega^{\circ}\cup\Omega_{B}$, this being the set of matrices characterized by the inequalities $a_{1},a_{2},b_{1},b_{2},c_{1},c_{2},N(u)\neq 0$.
Proposition 5.1.
The subset
$X=\{(u,v)\in\Omega^{\circ}\times\Omega^{\circ}|\Delta(u)=\Delta(v^{\ast})\}$
is irreducible in the Zariski topology.
Proof.
Let $\alpha=(\alpha_{1},\alpha_{2})\in(\mathbb{C}^{\times})^{2}$ and let $L_{\alpha}=\{u\in\Omega^{\circ}|\Delta(u)=\alpha\}$, $R_{\alpha}=\{v\in\Omega^{\circ}|\Delta(v^{\ast})=\alpha\}$. Then $X$ is the union of the fibers of the map $\varphi:X\longrightarrow\mathbb{C}^{\times}$ mapping $(u,v)\in X$ to $\Delta(u)=\Delta(v^{\ast})$. Now consider $L_{\alpha}$. This is the locus of
$\begin{array}[]{c}a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)-\alpha_{1}a_{1}(u)b_{1}(u)=0\\ a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)-\alpha_{2}a_{2}(u)b_{1}(u)=0.\end{array}$
It is the intersection of two quadrics and as such it is irreducible if $\alpha$ is in general position. Similarly $R_{\alpha}$ is also irreducible. Thus the fiber $\varphi^{-1}(\alpha)\cong L_{\alpha}\times R_{\alpha}$ is irreducible for $\alpha$ in general position. Thus $X$ is fibered over (×$(\mathbb{C}^{\times})^{2}$ with generically irreducible fibers, and it is therefore irreducible. ∎
Proposition 5.2.
Suppose $u,v$ are elements of $\Omega^{\circ}$ such that $\Delta(u)=\Delta(v^{\ast})$. Let $w$ be the normalized solution to the $\llbracket u,w,v\rrbracket=0$ as in Theorem 3.1. then
(17) $a_{1}(v)b_{1}(v)N(w)=a_{1}(w)b_{1}(w)N(v),\qquad a_{2}(v)b_{2}(v)N(w)=a_{2}(w)b_{2}(w)N(v)$
and
(18) $a_{1}(w)b_{2}(w)N(u)=a_{1}(u)b_{2}(u)N(w),\qquad a_{2}(w)b_{1}(w)N(u)=a_{2}(u)b_{1}(u)N(w),$
Remark 3.
Beginning with this Proposition, we will make use of continuity arguments in the sequel. This is for convenience and efficiency and could be replaced by calculations with equations. To emphasize this point, we will give two proofs of this result, one using a continuity argument, and one arguing directly from the equations.
First Proof.
First assume that $w\in\Omega^{\circ}$. Then these identities follow from Proposition 3.4 and equations (3), (4) and (5). We will deduce the general case of (17) from this special case by continuity. Let $X$ be as in Proposition 5.1. The subset of $X$ where $w\in\Omega^{\circ}$ is open in $X$. But $X$ is irreducible, so every nonempty open set is dense. Thus (17) and (18) are true on all of $X$ by continuity. ∎
Second Proof.
Alternative to the continuity argument in the first proof, we may modify the reasoning in the proof of Theorem 3.1 and deduce a formula for $u$ in terms of $w$ and $v$. It is important that the argument is valid without assuming that $w\in\Omega^{\circ}$. We rewrite the normalization condition (8) as
$c_{1}(u)=c_{1}(w)/c_{1}(v),\quad c_{2}(u)=c_{2}(w)/c_{1}(v).$
Substituting this into the equations (12), clearing the denominators and eliminating the redundant equations, we obtain
$\begin{array}[]{rcc}b_{2}(u)c_{1}(v)c_{2}(v)-b_{2}(w)a_{1}(v)+a_{1}(w)b_{2}(v)&=&0\\ -a_{1}(u)a_{1}(v)+b_{2}(u)b_{1}(v)+a_{1}(w)&=&0\\ -b_{1}(u)a_{1}(w)c_{1}(v)c_{2}(v)+a_{1}(u)b_{1}(w)c_{1}(v)c_{2}(v)-c_{1}(w)c_{2}(w)b_{1}(v)&=&0\\ -b_{2}(u)a_{2}(w)c_{1}(v)c_{2}(v)+a_{2}(u)b_{2}(w)c_{1}(v)c_{2}(v)-c_{1}(w)c_{2}(w)b_{2}(v)&=&0\\ -a_{2}(u)a_{2}(v)+b_{1}(u)b_{2}(v)+a_{2}(w)&=&0\\ b_{1}(u)c_{1}(v)c_{2}(v)-b_{1}(w)a_{2}(v)+a_{2}(w)b_{1}(v)&=&0\end{array}$
Since $v\in\Omega^{\circ}$, $a_{1}(v)$ and $a_{2}(v)$ are nonzero so
$a_{1}(u)=\frac{b_{2}(u)b_{1}(v)+a_{1}(w)}{a_{1}(v)},\qquad a_{2}(u)=\frac{b_{1}(u)b_{2}(v)+a_{2}(w)}{a_{2}(v)}.$
Substituting these there are eight equations but only four nonredundant ones. Among these are two that may be rearranged as:
$b_{1}(u)=\frac{b_{1}(w)a_{2}(v)-a_{2}(w)b_{1}(v)}{c_{1}(v)c_{2}(v)},\qquad b_{2}(u)=\frac{b_{2}(w)a_{1}(v)-a_{1}(w)b_{2}(v)}{c_{1}(v)c_{2}(v)}.$
Substituting these values, only two nonredundant equations remain, and these are (17). Equations (18) may be proved the same way, solving for $v$ instead of $u.$
Lemma 5.3.
Suppose that $u,v$ are elements of $\Omega^{\circ}$ such that $\Delta(u)=\Delta(v^{\ast})$. Let $w$ be the normalized solution to the $\llbracket u,w,v\rrbracket=0$ as in Theorem 3.1. Assume that $N(w)=0$. Then either $w\in\Omega_{a}$ or $w\in\Omega_{b}$.
Proof.
Since $u,v\in\Omega^{\circ}$, we have $N(u).N(v)\neq 0$. Since $N(w)=0$, equations (17) and (18) imply that $a_{1}(w)b_{1}(w)$, $a_{1}(w)b_{2}(w)$, $a_{2}(w)b_{1}(w)$ and $a_{2}(w)b_{2}(w)=0$. Therefore either $a_{1}(w)=a_{2}(w)=0$ or $b_{1}(w)=b_{2}(w)=0$. Suppose that $b_{1}(w)=b_{2}(w)=0$. Then $N(w)=a_{1}(w)a_{2}(w)-c_{1}(w)c_{2}(w)$ and since $N(w)$ and $c_{1}(w)c_{2}(w)$ both vanish, we must have $a_{1}(w)$ and $a_{2}(w)$ both nonzero and $w\in\Omega_{b}$. In the other case where $a_{1}(w)=a_{2}(w)=0$ similar reasoning shows that $w\in\Omega_{b}$. ∎
Proposition 5.4.
If $u,v\in\Omega^{\circ}$ such that $\Delta(u)=\Delta(v^{\ast})$, and if $w$ is the normalized solution to $\llbracket u,w,v\rrbracket=0$ guaranteed by Theorem 3.1. Then $w\in\overline{\Omega}$.
Proof.
We must prove that $w$ is in one of the sets $\Omega^{\circ}$, $\Omega_{b}$, $\Omega_{a}$ or $\Omega_{B}$. We may assume that $w\notin\Omega^{\circ}$ so at least one of $a_{1}(w)$, $a_{2}(w)$, $b_{1}(w)$, $b_{2}(w)$, $c_{1}(w)c_{2}(w)-b_{1}(w)b_{2}(w)$ or $N(w)$ is zero. The cases where $N(w)=0$ are handled by Lemma 5.3. Therefore we assume that $N(w)\neq 0$. The left side of (17) and (18) is then nonzero and so these equations imply that $a_{1}(w),a_{2}(w),b_{1}(w)$ and $b_{2}(w)$ are all nonzero. Therefore $c_{1}(w)c_{2}(w)-b_{1}(w)b_{2}(w)=0$ and $w\in\Omega_{B}$. ∎
Proposition 5.5.
Suppose that $u,v\in\overline{\Omega}$ and at least one of $u,v\in\Omega_{b}\cup\Omega_{a}$. Then there exists $w\in\Omega$ such that $\llbracket u,w,v\rrbracket=0$. The solution is unique unless both $u,v\in\Omega_{a}$.
Remark 4.
We assume that $u\in\Omega_{b}\cup\Omega_{a}$, leaving the other case to the reader. It may be surpising that a solution always exists, in contrast with Theorem 3.1. But note that $\Delta(u)$ is not defined if $u\in\Omega_{a}$ or $\Omega_{b}$ since the numerator and denominator in (3) and (4) both vanish. Therefore the condition $\Delta(u)=\Delta(v^{\ast})$ has no meaning. This makes it less surprising that the solution exists without condition.
Proof.
If $u\in\Omega_{b}$, then solving the Yang-Baxter equation as in the proof of Theorem 3.1 leads to the unique normalized solution
$\begin{array}[]{lll}a_{1}(w)=a_{1}(u)a_{1}(v),&&a_{2}(w)=a_{2}(u)a_{2}(v),\\ b_{1}(w)=a_{2}(u)b_{1}(v),&&b_{2}(w)=a_{1}(u)b_{2}(v).\\ c_{1}(w)=c_{1}(u)c_{1}(v).&&c_{2}(w)=c_{2}(u)c_{2}(v).\end{array}$
With these values, there is only one remaining equation $a_{1}(u)a_{2}(u)=c_{1}(u)c_{2}(u)$, which is automatic since $u\in\Omega_{b}$.
If $u\in\Omega_{a}$, then we obtain, we obtain
$\begin{array}[]{lll}a_{1}(w)=-b_{2}(u)b_{1}(v),&&a_{2}(w)=-b_{1}(u)a_{2}(v),\\ c_{1}(w)=c_{1}(u)c_{1}(v).&&c_{2}(w)=c_{2}(u)c_{2}(v).\end{array}$
Substituting these values, three equations remain. One is the condition $b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)=0$, which is automatically satisfied since $u\in\Omega_{a}$. The other two equations are
$\begin{array}[]{l}b_{2}(u)(c_{1}(v)c_{2}(v)-b_{1}(v)b_{2}(v))=b_{2}(w)a_{1}(v),\\ b_{1}(u)(c_{1}(v)c_{2}(v)-b_{1}(v)b_{2}(v))=b_{1}(w)a_{2}(v).\end{array}$
If $v\notin\Omega_{a}$, these conditions determine $b_{1}(w)$ and $b_{2}(w)$ uniquely. On the other hand if $v\in\Omega_{a}$ then both sides vanish, so $b_{1}(w)$ and $b_{2}(w)$ are unconstrained. ∎
6. The Six-Vertex Yang-Baxter Groupoid
Roughly, the idea is to define a composition law on six-vertex matrices by writing $w=u\star v$ if $\llbracket u,w,v\rrbracket=0$. Usually this determines $w$ up to a constant, which we can fix by requiring the solution to be normalized. For the groupoid inverse, which we will denote by $u^{\prime}$, we can take an appropriate multiple of the matrix inverse $u^{-1}$. More precisely, we find $u^{\ast}$ to be more convenient to work with, and we define $u^{\prime}=\frac{1}{c_{1}(u)c_{2}(u)}u^{\ast}$. These definitions must be modified in special cases: for example $u$ may not be invertible but we still want $u^{\prime}$ to be defined.
The above description is only approximately correct. We can begin by dividing the groupoid into the disjoint union of two parts, the free-fermionic groupoid, and the non-free-fermionic groupoid. These may be handled separately. Actually the free-fermionic part is a group, isomorphic to $\operatorname{GL}(2,\mathbb{C})\times\operatorname{GL}(1,\mathbb{C})$, and we have already treated it in Section 4. Thus it remains to construct the non-free-fermionic groupoid $\mathfrak{G}_{\operatorname{nf}}$. As a set, we may start with $\overline{\Omega}=\Omega^{\circ}\cup\Omega_{b}\cup\Omega_{a}\cup\Omega_{B}$. Unfortunately, we need $\Delta_{1},\Delta_{2}$ and $u^{\ast}$ to be defined on the whole groupoid, but $\Delta$ is undefined on $\Omega_{b}$ and $\Omega_{a}$, because the numerator and denominator in (3) and (4) both vanish. Moreover $u^{\ast}$ is undefined on $\Omega_{a}$ because the numerator and denominator in (3) both vanish.
To fix these problems, we “blow up” $\Omega_{b}$ and $\Omega_{a}$. We define $\mathfrak{G}_{\operatorname{nf}}$ as a set to be
$\left\{(u,d_{1},d_{2})\in\overline{\Omega}\times\mathbb{C}^{\times}\times\mathbb{C}^{\times}|a_{1}(u)b_{1}(u)d_{1}=a_{2}(u)b_{2}(u)d_{2}=N(u)\right\}.$
The map $\pi:\mathfrak{G}_{\operatorname{nf}}\longrightarrow\overline{\Omega}$ is the projection on the first component. Note that over $\Omega=\Omega^{\circ}\cup\Omega_{B}$, the fibers of $\pi$ have cardinality 1. That is, there is a unique section $\mathbf{s}:\Omega\longrightarrow\mathfrak{G}_{\operatorname{nf}}$ that maps $u\in\Omega$ to $\mathbf{s}(u)=(u,\Delta_{1}(u),\Delta_{2}(u))$ and $\pi^{-1}(u)=\{\mathbf{s}(u)\}$.
On the other hand, suppose that $u$ is in either $\Omega_{a}$ or $\Omega_{b}$. Then $\pi^{-1}(u)$ consists of a 2-dimensional torus $\{(u,d_{1},d_{2})|d_{i}\in\mathbb{C}^{\times}\}$. This is because $a_{1}(u)b_{1}(u)$, $a_{2}(u)b_{2}(u)$ and $N(u)$ are all zero, so there is no constraint on $d_{1}$ and $d_{2}$.
We will adopt the following convention. If $\mathbf{u}\in\mathfrak{G}_{\operatorname{nf}}$ we will use the same letter $u$ to denote the element $\pi(\mathbf{u})\in\overline{\Omega}$. If $u\in\Omega=\Omega^{\circ}\cup\Omega_{B}$, we may even conflate $u\in\Omega$ with $\mathbf{s}(u)$. Thus we are identifying $\Omega$ with its corresponding subset of $\mathfrak{G}_{\operatorname{nf}}$. Therefore we consider $\Omega^{\circ}$ and $\Omega_{B}$ to be subsets of $\mathfrak{G}_{\operatorname{nf}}$.
The space $\Omega$ is 6-dimensional, and $\Omega_{B}$ is 5-dimensional. On the other hand, $\Omega_{b}$ and $\Omega_{a}$ are both 3-dimensional. But we define $\Gamma_{b}=\pi^{-1}(\Omega_{b})$ and $\Gamma_{a}=\pi^{-1}(\Omega_{a})$ and these are 5-dimensional. So all three “boundary components” $\Gamma_{b}$, $\Gamma_{a}$ and $\Omega_{B}$ are of codimension 1. We will see in Proposition 6.1 that the involution interchanges $\Gamma_{a}$ and $\Omega_{B}$, so even though their constructions are quite different, they are in that sense equivalent.
Our general strategy is to prove things generically for elements of the dense open set $\Omega^{\circ}$, then deduce them by continuity in general. We have already seen an example of this in the first proof of Proposition 5.2. Now $\mathfrak{G}_{\operatorname{nf}}$ is naturally a quasi-affine algebraic variety, irreducible by Proposition 5.1, with $\Omega^{\circ}$ a dense open set. But in this section we will switch to the complex topology for our continuity arguments. So let us say what it means for a sequence $\{r_{n}\}\subset\Omega^{\circ}$ to converge to an element $\mathbf{r}=(r,d_{1},d_{2})$, possibly of one of the boundary components $\Omega_{b}$ or $\Omega_{a}$. It means that $r_{n}\rightarrow r$ in the complex topology, and moreover $\Delta_{1}(r_{n})\rightarrow d_{1}$ and $\Delta_{2}(r_{n})\rightarrow d_{2}$. It is not hard to see that $\Omega^{\circ}$ is dense in $\mathfrak{G}_{\operatorname{nf}}$.
Proposition 6.1.
The map $r\mapsto r^{\ast}$ can be extended uniquely to a continuous map $\mathfrak{G}_{\operatorname{nf}}\longrightarrow\mathfrak{G}_{\operatorname{nf}}$. The map preserves $\Omega^{\circ}$ and $\Gamma_{b}$ but interchanges $\Gamma_{a}$ and $\Omega_{B}$. If $\mathbf{r}=(r,d_{1},d_{2})$ then $\mathbf{r}^{\ast}=(r^{\ast},d_{1}^{\ast},d_{2}^{\ast})$ where
(19) $r^{\ast}=\left(\begin{array}[]{cccc}a_{1}(\mathbf{r}^{\ast})&&&\\ &c_{2}(r)&-b_{1}(r)&\\ &-b_{2}(r)&c_{1}(r)&\\ &&&a_{2}(\mathbf{r}^{\ast})\end{array}\right)$
with
$(a_{1}(\mathbf{r}^{\ast}),a_{2}(\mathbf{r}^{\ast}))=\left\{\begin{array}[]{ll}(a_{2}(r),a_{1}(r))&\text{if }\mathbf{r}\in\Gamma_{b},\\ (-d_{1}b_{1}(r),-d_{2}b_{2}(r))&\text{if $\mathbf{r}\in\Gamma_{a}$,}\\ (0,0)&\text{if }\mathbf{r}\in\Omega_{B}\end{array}\right.$
and
$(D_{1}^{\ast},D_{2}^{\ast})=\left\{\begin{array}[]{ll}\left(\frac{b_{2}(r)}{b_{1}(r)}d_{2},\frac{b_{1}(r)}{b_{2}(r)}d_{1}\right)&\text{if $\mathbf{r}\in\Gamma_{a}$,}\\ \left(\frac{a_{1}(r)}{a_{2}(r)}d_{1},\frac{a_{2}(r)}{a_{1}(r)}d_{2}\right)&\text{otherwise.}\end{array}\right.$
Proof.
We must consider the extension to $\Gamma_{b}$, $\Gamma_{a}$ and $\Omega_{B}$. We consider $\mathbf{r}=(r,d_{1},d_{2})$ in one of these boundary components, and a sequence $\{r_{n}\}\subset\Omega^{\circ}$ that converges to $\mathbf{r}$.
If $r\in\Omega_{B}$ then $a_{1}^{\ast}$ and $a_{2}^{\ast}$ are continuous at $r$ and converge to zero. For the other two cases, note that
$a_{1}^{\ast}(r_{n})=\frac{a_{1}(r_{n})a_{2}(r_{n})-N(r_{n})}{a_{1}(r_{n})}=a_{2}(r_{n})-b_{1}(r_{n})\Delta_{1}(r_{n}).$
Thus
$a_{1}^{\ast}(r_{n})\longrightarrow\left\{\begin{array}[]{ll}a_{2}(r)&\text{if $r\in\Omega_{b}$,}\\ -D_{1}b_{1}(r)&\text{if }r\in\Omega_{a}.\\ 0&\text{if }r\in\Omega_{B},\end{array}\right.\qquad a_{2}^{\ast}(r_{n})\longrightarrow\left\{\begin{array}[]{ll}a_{1}(r)&\text{if $r\in\Omega_{b}$,}\\ -D_{2}b_{2}(r)&\text{if }r\in\Omega_{a}.\\ 0&\text{if }r\in\Omega_{B},\end{array}\right.$
We must also consider the limits of the $\Delta_{i}(r_{n}^{\ast})$. If $r\in\Omega_{b}$ or $\Omega_{B}$ then
$\Delta_{1}(r_{n}^{\ast})=\frac{a_{1}(r_{n})}{a_{2}(r_{n})}\Delta_{1}(r_{n})\longrightarrow\frac{a_{1}(r)}{a_{2}(r)}D_{1}.$
Let us assume that $r\in\Omega_{a}$. Then $r_{n}^{\ast}\rightarrow r^{\ast}$ where $r^{\ast}$ is already computed. We have $N(r^{\ast})=a_{1}(r^{\ast})a_{2}(r^{\ast})-\det(B(r^{\ast}))=a_{1}(r^{\ast})a_{2}(r^{\ast})$ so $\Delta_{1}(r^{\ast})=a_{2}(r^{\ast})/(-b_{1}(r))=D_{2}b_{2}(r)/b_{1}(r)$. ∎
Corollary 6.2.
We have
$\Delta(\mathbf{r}^{\ast})=\left\{\begin{array}[]{ll}\left(\frac{a_{1}(r)}{a_{2}(r)}\Delta_{1}(\mathbf{r}),\frac{a_{1}(r)}{a_{2}(r)}\Delta_{2}(\mathbf{r})\right)&\text{if $\mathbf{r}\notin\Gamma_{a}$},\\ \left(\frac{b_{2}(r)}{b_{1}(r)}\Delta_{2}(\mathbf{r}),\frac{b_{1}(r)}{b_{2}(r)}\Delta_{1}(\mathbf{r})\right)&\operatorname{if}\mathbf{r}\in\Gamma_{a}.\end{array}\right.$
Corollary 6.3.
The map $r\mapsto r^{\prime}$, where $r^{\prime}=\frac{1}{c_{1}(r)c_{2}(r)}r^{\ast}$ extends to a continuous map $\mathbf{r}\rightarrow\mathbf{r}^{\prime}$ of $\mathfrak{G}_{\operatorname{nf}}$.
Proof.
Since $c_{1}(\mathbf{r})$ anc $c_{2}(\mathbf{r})$ are defined on the entire groupoid. ∎
Lemma 6.4.
Let $\mathbf{r}\in\mathfrak{G}_{\operatorname{nf}}$. Then there is a sequence $\{r_{n}\}\subset\Omega^{\circ}$ such that $r_{n}\rightarrow\mathbf{r}$, with $\Delta(r_{n})=\Delta(\mathbf{r})$ and $\Delta(r_{n}^{\ast})=\Delta(\mathbf{r})$.
Proof.
If $\mathbf{r}=(r,d_{1},d_{2})\in\Omega^{\circ}$ we may take $r_{n}=r$, so this case is obvious. We need to check this if $\mathbf{r}$ is in one of the boundary components.
First suppose $\mathbf{r}\in\Gamma_{b}$, so $b_{1}(r)=b_{2}(r)=0$ and $a_{1}(r)a_{2}(r)=c_{1}(r)c_{2}(r)$. We define the sequence $r_{n}$ so that $a_{1}(r_{n})=a_{1}(r)$ and $a_{2}(r_{n})=a_{2}(r)$. Let $(d_{1},d_{2})=\Delta(\mathbf{r})$. In order that $\Delta(r_{n})=(d_{1},d_{2})$ we need:
(20) $\frac{N(r_{n})}{a_{1}(r)b_{1}(r_{n})}=d_{1},\qquad\frac{N(r_{n})}{a_{2}(r)b_{2}(r_{n})}=d_{2}.$
Thus we want $b_{1}(r_{n}),b_{2}(r_{n})\rightarrow 0$ preserving the ratio
$\frac{b_{1}(r_{n})}{b_{2}(r_{n})}=\frac{d_{2}a_{2}(r)}{d_{1}a_{1}(r)}.$
Then if the first equation in (20) is satisfied, the second is automatically true.
If $c$ denotes the constant on the right, then we can choose a sequence $b_{2}(r_{n})\rightarrow 0$ and define $b_{1}(r_{n})=c\cdot b_{2}(r_{n})$. Now we need:
$a_{1}(r)a_{2}(r)+b_{1}(r_{n})b_{2}(r_{n})-c_{1}(r_{n})c_{2}(r_{n})=d_{1}a_{1}(r)b_{1}(r_{n}).$
We can fix $c_{1}(r_{n})=c_{1}(r)$ and then solve this equation for $c_{2}(r_{n})$. The right-hand side converges to zero as does the second term on the left, so
$a_{1}(r)a_{2}(r)-c_{1}(r)c_{2}(r_{n})\rightarrow 0.$
Since $a_{1}(r)a_{2}(r)=c_{1}(r)c_{2}(r)$ it follows that $c_{2}(r_{n})\rightarrow c_{2}(r)$.
We also need to know that $\Delta(r_{n}^{\ast})=\Delta(\mathbf{r}^{\ast})$. Indeed
$\Delta_{1}(\mathbf{r}^{\ast})=\frac{a_{1}(r)}{a_{2}(r)}\Delta_{1}(\mathbf{r})=\frac{a_{1}(r_{n})}{a_{2}(r_{n})}\Delta_{1}(r_{n})=\Delta_{1}(r_{n}^{\ast})$
and similarly for $\Delta_{2}$.
We leave the cases where $\mathbf{r}\in\Gamma_{a}$ or $\Omega_{B}$ to the reader, except to note that the cases are equivalent since if $r_{n}\rightarrow\mathbf{r}$ then $r_{n}^{\ast}\rightarrow\mathbf{r}^{\ast}$ with $\mathbf{r}\in\Gamma_{a}$ and $\mathbf{r}^{\ast}\in\Omega_{B}$, so one sequence works for both cases. ∎
Now let us define the groupoid composition $\star$ on $\mathfrak{G}_{\operatorname{nf}}$.
Proposition 6.5.
Let $\mathbf{u},\mathbf{v}\in\mathfrak{G}_{\operatorname{nf}}$. Assume that $\Delta(\mathbf{u})=\Delta(\mathbf{v}^{\star})$. Then there is a unique element $\mathbf{w}$ such that $\llbracket u,w,v\rrbracket=0$ is a normalized solution, and such that $\Delta(\mathbf{w})=\Delta(\mathbf{v})$ and $\Delta(\mathbf{w}^{\ast})=\Delta(\mathbf{u}^{\ast})$. The element $\mathbf{w}$ depends continuously on $\mathbf{u}$ and $\mathbf{v}$.
Proof.
This follows from Theorem 3.1 if $\mathbf{u},\mathbf{v}\in\Omega$, but we must consider the case where one or both is in $\Omega_{b}$ or $\Omega_{a}$. In this case, the existence of $w$ such that $\llbracket u,w,v\rrbracket=0$ is guaranteed by Proposition 5.5. However if both $\mathbf{u},\mathbf{v}\in\Gamma_{a}$ that Proposition shows that while $a_{1}(\mathbf{w}),a_{2}(\mathbf{w}),c_{1}(\mathbf{w}),c_{2}(\mathbf{w})$ are determined, but we must deduce the values for $b_{1}(\mathbf{w})$ and $b_{2}(\mathbf{w})$ from the requirement that $\Delta(\mathbf{w})=\Delta(\mathbf{v})$. We consider the identities
$b_{1}(r)=\frac{a_{2}(r)-a_{1}(r^{\ast})}{\Delta_{1}(r)},\qquad b_{1}(r)=\frac{a_{1}(r)-a_{2}(r^{\ast})}{\Delta_{2}(r)},$
which follow from the definitions of $\Delta_{i}$ and $r^{\ast}$ for $r\in\Omega$, and by continuity to $\mathfrak{G}_{\operatorname{nr}}$. Since we require $\Delta(\mathbf{w})=\Delta(\mathbf{v})$, we see that we must define
$b_{1}(\mathbf{w})=\frac{a_{2}(\mathbf{w})-a_{1}(\mathbf{w}^{\ast})}{\Delta_{1}(\mathbf{v})},\qquad b_{2}(\mathbf{w})=\frac{a_{1}(\mathbf{w})-a_{2}(\mathbf{w}^{\ast})}{\Delta_{2}(\mathbf{v})}.$
The continuity of the $\star$ operation is clear. ∎
If $\Delta(\mathbf{u})=\Delta(\mathbf{v}^{\ast})$, let $\mathbf{u}\star\mathbf{v}=\mathbf{w}$, where $\mathbf{w}$ is as in Proposition 6.5. In this case we say that $\mathbf{u}\star\mathbf{v}$ is defined
Proposition 6.6 (Associativity for general position).
If $r,s,t,u,v\in\Omega^{\circ}$ are such that $r\star t=s$ and $t\star v=u$. Then $s\star v$ and $r\star u$ are both defined, and they are equal.
Proof.
Note that although by assumption $r,s,t,u,v$ are all in $\Omega^{\circ}$ the compositions $s\star v$ and $r\star u$ could be in a boundary component.
By Proposition 3.4, $\Delta(r)=\Delta(t^{\ast})$ and $\Delta(s)=\Delta(t)$, and similarly $\Delta(t)=\Delta(v^{\ast})$, and $\Delta(u)=\Delta(v)$. Note that $\Delta(s)=\Delta(t)=\Delta(v^{\ast})$ implies that $s\star v$ is defined by Proposition 6.5. On the other hand, we note that $\llbracket t,u,v\rrbracket=0$ since $u=t\star v$, we also have $\llbracket v,t^{\ast},u^{\ast}\rrbracket=0$. Thus $\Delta(t^{\ast})=\Delta(u^{\ast})$. Therefore $\Delta(r)=\Delta(u^{\ast})$ and so $r\star u$ is defined.
It remains to be shown that $r\star u=s\star v$. We have
$c_{1}(r\star u)=c_{1}(r)c_{1}(t)\star c_{1}(v)=c_{1}(s\ast v),$
and similarly for $c_{2}$. Using (10) we have
$\begin{array}[]{lll}a_{1}(s\star v)&=&a_{1}(s)a_{1}(v)-b_{2}(s)b_{1}(v)\\ &=&(a_{1}(r)a_{1}(t)-b_{2}(r)b_{1}(t))a_{1}(v)-(a_{1}(r)b_{2}(t)+b_{2}(r)a_{1}^{\ast}(t))b_{1}(v)\\ &=&a_{1}(r)(a_{1}(t)a_{1}(v)-b_{2}(t)b_{1}(v))-b_{2}(r)(b_{1}(t)a_{1}(v)+a_{1}(t^{\ast})b_{1}(v))\\ &=&a_{1}(r)a_{1}(u)-b_{2}(r)b_{1}(u)=a_{1}(r\star u)\end{array}$
and similarly $a_{2}(s\star v)=a_{2}(r\star u)$.
We note that $s^{\ast}=t^{\ast}\star r^{\ast}$ by Lemma 3.3. Therefore using (10) and the fact that $b_{1}(r^{\ast})=-b_{1}(r)$ and $b_{2}(t^{\ast})=-b_{2}(t)$ we have
$a_{1}(s^{\ast})=a_{1}(r^{\ast})a_{1}(t^{\ast})-b_{1}(r^{\ast})b_{2}(t^{\ast})=a_{1}(r^{\ast})a_{1}(t^{\ast})-b_{1}(r)b_{2}(\Delta t).$
Using these equations
$\begin{array}[]{lll}b_{1}(s\star v)&=&a_{1}(s^{\ast})b_{1}(v)+b_{1}(s)a_{1}(v)\\ &=&(a_{1}(t^{\ast})a_{1}(r^{\ast})-b_{1}(r)b_{2}(t))b_{1}(v)+(b_{1}(r)a_{1}(t)+b_{1}(t)a_{1}(r^{\ast}))a_{1}(v),\\ &&\Delta\\ b_{1}(r\star u)&=&a_{1}(r^{\ast})b_{1}(u)+b_{1}(r)a_{1}(u)\\ &=&a_{1}(r^{\ast})(a_{1}(t^{\ast})b_{1}(v)+b_{1}(t)a_{1}(v))+b_{1}(r)(a_{1}(t)a_{1}(v)-b_{2}(t)b_{1}(v)).\end{array}$
These are term by term equal, and similarly $b_{2}(s\star v)=b_{2}(r\star u)$. We see that $s\star v$ and $r\star u$ have the same Boltzmann weights, and so they are equal. ∎
The next two results establish that the composition law that we have defined on $\mathfrak{G}_{\operatorname{nf}}$ satisfies the groupoid axioms.
Theorem 6.7 (Associativity Axiom).
Let $\mathbf{r},\mathbf{s},\mathbf{t},\mathbf{u},\mathbf{v}\in\mathfrak{G}_{\operatorname{nf}}$ such that $\mathbf{r}\star\mathbf{t}=\mathbf{s}$ and $\mathbf{t}\star\mathbf{v}=\mathbf{u}$. Then  are defined, and they are equal.
Proof.
We could argue similarly to Proposition 6.6, on a case-by-case basis to handle the edge cases, where one or more of $\mathbf{r},\mathbf{s},\mathbf{t},\mathbf{u},\mathbf{v}$ is in one of the boundary components. Instead, we will argue by continuity. By Lemma 6.4 we may find sequences $\{r_{n}\},\{t_{n}\},\{v_{n}\}\subset\Omega^{\circ}$ such that $\Delta(r_{n})=\Delta(\mathbf{r})$ with $r_{n}\rightarrow\mathbf{r}$, etc. Then for each $n$, $\Delta(r_{n})=\Delta(t_{n}^{\ast})$ so $r_{n}\star t_{n}$ is defined; call this $s_{n}$. With $\{t_{n}\}$ fixed, we may perturb the sequence $\{r_{n}\}$ so that $s_{n}\in\Omega^{\circ}$. Similarly, with $\{t_{n}\}$ fixed, we may purturb the sequence $\{v_{n}\}$ so that $u_{n}\in\Omega^{\circ}$. Now by Proposition 6.6, $r_{n}\star u_{n}$ and $s_{n}\star v_{n}$ are defined and equal, and taking the limit gives $\mathbf{s}\star\mathbf{v}=\mathbf{r}\star\mathbf{u}$. ∎
If $\mathbf{u}$, we define the groupoid inverse to be $\mathbf{u}^{\prime}:=\frac{1}{c_{1}(u)c_{2}(u)}\mathbf{u}^{\ast}$.
Proposition 6.8 (Idempotents).
Let $d_{1},d_{2}\in\mathbb{C}^{\times}$. Let $\mathbf{I}_{d_{1},d_{2}}=(I_{V\otimes V},d_{1},d_{2})$. Then $\mathbf{I}_{d_{1},d_{2}}$ is an idempotent in that $\mathbf{I}_{d_{1},d_{2}}\star\mathbf{I}_{d_{1},d_{2}}=\mathbf{I}_{d_{1},d_{2}}$. We have $\mathbf{I}_{d_{1},d_{2}}^{\ast}=\mathbf{I}_{d_{1},d_{2}}^{\prime}=\mathbf{I}_{d_{1},d_{2}}$. Furthermore $\Delta(\mathbf{I}_{d_{1},d_{2}})=(d_{1},d_{2})$. If $\mathbf{u}\star\mathbf{I}_{d_{1},d_{2}}$ is defined $\mathbf{u}\star\mathbf{I}_{d_{1},d_{2}}=\mathbf{u}$. If $\mathbf{I}_{d_{1},d_{2}}\star\mathbf{v}$ is defined then $\mathbf{I}_{d_{1},d_{2}}\star\mathbf{v}=\mathbf{v}$.
Proof.
Proposition 6.1 shows that $\mathbf{I}_{d_{1},d_{2}}^{\ast}=\mathbf{I}_{d_{1},d_{2}}$. Therefore $\mathbf{I}_{d_{1},d_{2}}\star\mathbf{I}_{d_{1},d_{2}}$ is defined. It is of the form $\mathbf{i}=(J,D_{1},D_{2})$, where by the uniqueness assertion in Proposition 5.5, $J$ is the unique element of $\overline{\Omega}$ such that $\llbracket I,J,I\rrbracket=0$ is a normalized solution; thus $J=I$, and by Proposition 6.5 we have $(D_{1},D_{2})=\Delta(\mathbf{i})=\Delta(I_{d_{1},d_{2}})=(d_{1},d_{2})$. Therefore $\mathbf{i}=\mathbf{I}_{d_{1},d_{2}}$. We leave the remaining details to the reader. ∎
Theorem 6.9 (Inverse Axiom).
Let $\mathbf{u}\in\mathfrak{G}_{\operatorname{nf}}$. Then $\mathbf{u}\star\mathbf{u}^{\prime}=\mathbf{I}_{\Delta(\mathbf{u}^{\prime})}$ and $\mathbf{u}^{\prime}\star\mathbf{u}=\mathbf{I}_{\Delta(u)}$ are defined. If $\mathbf{r}\star\mathbf{u}$ is defined then $(\mathbf{r}\star\mathbf{u})\star\mathbf{u}^{\prime}=\mathbf{r}$, while if $\mathbf{u}\star\mathbf{t}$ is defined then $\mathbf{u}^{\prime}\star(\mathbf{u}\star\mathbf{t})=\mathbf{t}$.
Proof.
First suppose that $u\in\Omega^{\circ}$. Then $u$ is invertible and
$u^{\prime}=\gamma(u)u^{-1},\qquad\gamma(u):=\frac{\det(B(u))}{c_{1}(u)c_{2}(u)}.$
We have (denoting $I_{V}$)
$\llbracket u,I_{V\otimes V},u^{-1}\rrbracket=(u\otimes I)(I\otimes I\otimes I)(u^{-1}\otimes I)-(I\otimes u^{-1})(I\otimes I\otimes I)(I\otimes u)=0,$
so $u\star u^{-1}$ is defined and is a constant multiple of $I_{V\otimes V}$. Now $u^{\prime}$ is a constant multiple of $u^{-1}$, so $u\star u^{\prime}$ is a constant multiple of $I_{V\otimes V}$ and it may be checked that $\llbracket u,I,u^{\prime}\rrbracket=0$ is the normalized solution, so indeed $\mathbf{u}\star\mathbf{u}^{\prime}=I_{d_{1},d_{2}}$ where $(d_{1},d_{2})=\Delta(\mathbf{u}^{\prime})=\Delta(\mathbf{u}^{\ast})$. The last assertion follows from associativity and Proposition 6.8.
We have assumed that $u\in\Omega^{\circ}$. For $\mathbf{u}$ in one of the boundary components, By Lemma 6.4 we may chose a sequence $\{u_{n}\}\subset\Omega^{\circ}$ that converges to $\mathbf{u}$ and such that $\Delta(u_{n})=\Delta(\mathbf{u})$. Then if $\mathbf{r}\star\mathbf{u}$ is defined, so is $\mathbf{r}\star u_{n}$, and we may deduce the general result by continuity. ∎
From the last three results, we see that $\mathfrak{G}_{\operatorname{nf}}$ is a groupoid. We will call the disjoint union $\mathfrak{G}_{\operatorname{ff}}\sqcup\mathfrak{G}_{\operatorname{nf}}$ the six-vertex groupoid, which accounts for all Yang-Baxter equations that we may construct from the six-vertex model.